# Variational autoencoders stabilise TCN performance when classifying weakly labelled bioacoustics data

### v1

@@ 3.2.1 Extraction of clips for training feature engineering methods
An impulsive noise detector algorithm (Algorithm 1) was used to identify and select up to 10 transient sounds with the highest amplitude within each 4 minute file, regardless of what the recording had been labeled as. Subsequently, waveform segments of varying lengths were extracted from these transients to train different feature engineering methods, ensuring the random placement of amplitude peaks within each segment. Additionally, randomly sampled segments where no transients was detected were included in order to incorporate examples of background noise into the optimisation of feature engineering methods. The proposed approach acknowledges that, while impulsive noises detected in a 4-min sequence might not necessarily share the label of the parent audio file, a considerable proportion are expected to do so, such that this procedure would produce a dataset representative of the different sound sources present in the data.
3.2.2 Handcrafted feature generation
@@
Several acoustic features were extracted from the non-overlapping short audio segments generated as described in Subsection 3.2.1, namely the root-mean-square of the time-series for a specific frequency band $[f_{1},f_{2}]$ ($RMS_{f_{1},f_{2}}$), peak and mean frequencies, spectral width, and energy sum over three frequency bands (1kHz - 4kHz, 4kHz-8kHz, 8kHz-16kHz) (see Figure 2 for an example and Table 2 for formal definition of each handcrafted feature). The choice of these parameters was motivated by earlier works on the detection-classification of odontocete vocalisations [12], and the frequency bands were chosen to correspond with distinct bandwidths in the sperm whale click spectral profile [71].
@@
Table 2: Handpicked acoustic features used in the study
@@
Peak frequency $pkf=f_{i}|\left(s\left[f_{i}\right]=max(s)\right)$ $f$: vector of frequencies returned by fft
@@ 5 $RMS_{1-2kHz}$, $RMS_{2-4kHz}$, $RMS_{4-8kHz}$, $RMS_{8-16kHz}$, $RMS_{16-20kHz}$,
Spectral features 4 $RMS_{1-20kHz}$, $pkf$, $\overline{f}$, $\omega$
7 $RMS_{1-20kHz}$, $pkf$, $\overline{f}$, $\omega$, $E_{1-4kHz}$, $E_{4-8kHz}$, $E_{8-16kHz}$
3.2.3 Feature Extraction with Variational AutoEncoders (VAE)
VAEs were trained on three different forms of input data: waveforms, spectral profiles, and spectrogram representations. To generate the spectrograms and spectral profiles we used a 512-point Hann window fft with a 50% overlap. Spectral profiles were computed on waveforms of length 512 samples, and frequencies below 1kHz or above 20kHz were discarded, leading to a final sequence of length 200, which was then smoothed using an 8-point moving average. Additionally, longer spectrograms were computed for waveform sections of length 32,768 samples (0.68 seconds), so as to have the resulting spectrogram be 128 FTTs in length giving an image of width 128 pixels. Again, frequencies below 1kHz and above 20kHz were discarded to center sperm whale clicks in the spectrogram [71], and the final representation was resized to be of size 128x128 pixels.
[…]
3.3 Training recording classifier
For audio classification at the file level, recordings were divided into shorter, non-overlapping windows of two sizes: 512 and 2048 samples (0.01s and 0.04s, respectively). The two window lengths were used to investigate the trade-off between temporal resolution and sequence length. This process generated 22,500 and 5,625 windows, respectively, for 4-minute files, and 2,812 and 703 windows for 30-second files (Table 4). The feature vectors detailed in the previous section were concatenated, transforming each audio file into a sequence of size ($m$, $n$), where $m$ represents the number of extracted features and $n$ denotes the number of non-overlapping windows in the file. Each of the $m$ features was standardised using median and standard deviations of the feature on the values obtained on the short audio segment dataset. This standarisation was done on both the handpicked features and the VAE embeddings.
A TCN was used for classifying full recordings from the sequence of extracted features based on the presence or absence of sperm whale vocalisations (Figure 3). This architecture was chosen due to is proven ability to capture long-term dependencies and timing patterns in acoustic signals [78], which would be key for detecting groups of sperm whales of varying sizes and distinguishing them from other sources of impulsive noise
The TCN was of length $n$ and had $m$ channels. Due to the range of different feature extraction methods tested, trained models varied in complexity (Table 4). All TCNs were trained in a supervised manner using the Adam optimizer and negative log likelihood loss function. Hyperparameters, including the number of hidden layers, kernel size, dropout rate, learning rate, and number of levels, were optimized using a grid search strategy (Table 7).
@@ 3.3 Training recording classifier
RMS 1, 3, 5 512 samples 22500 2812
2048 samples 5625 703
Spectral parameters 4, 7 512 samples 22500 2812
2048 samples 5625 703
1D-VAE on waveform 8, 16, 24, 32 512 samples 22500 2812
2048 samples 5625 703
1D-VAE on spectral profile 8, 16, 24, 32 512 samples 22500 2812
2D-VAE on spectrogram 8, 16, 24, 32, 64, 128 32768 samples 352 44
3.4 Classifier comparison
The performances metrics of the different trained TCNs were compared on a subset of the data reserved for validation. The training and validation split was random and consistent across models trained on the same length of recordings. Performance evaluation was conducted based on recall ($TP/P$) and false-positive-rate ($FPR$; $FP/N$), where $P$ and $N$ represent the total number of positive and negative files (as annotated), and $TP$ and $FN$ denote correct and incorrect predictions of the model from the files labelled as containing sperm whales by humans. The optimal model was determined by minimizing $FPR$ while simultaneously maximizing recall.
4 Results
4.1 Audio data annotation
The final 4-minute training set contained 9,352 audio files (Table 5) coming from 9 different deployments, with a 50/50 split between files that were deemed as containing and not containing sperm whale clicks. This dataset included data from different sperm whale populations spanning 20 years, 8 different geographical locations, 7 different recording devices and over 8 other vocalising species that produce echolocation clicks, as well as other sources of impulsive noise. The 30-second dataset contained a total of 14,138 files from 2 different sources, AS and MED (Table 5).
Table 5: The number of files in the merged dataset labelled as containing only sperm whales (SW), sperm whales and other sources of impulsive noise, other sources of impulsive noise only, and files audited but nothing noted down in the final training set separated by source. For the Balearic data files were only marked for the presence of sperm whale vocalisations, so no information on other sound sources was available. The numbers outside of parenthesis refer to the 4-minute dataset, and parenthesis to the 30-second dataset.
@@ 4.1 Audio data annotation
BAL_4 78 - - 78 156
CAL 175 1 116 60 352
CS 773 557 720 610 2660
ICE 365 558 379 544 1846
MED 617 (4067) 786 (1063) 58 (161) 1345 (4999) 2806 (10286)
Total 2766 (5963) 1910 (1089) 1411 (1216) 3265 (5870) 9352 (14138)
4.2 Clips extracted for training feature engineering methods
@@
A total of 162,679 short segments were identified to train the feature engineering methods (see as described in Subsection 3.2.1) to train the feature extraction methods. 112,869 of these were sampled using an impulsive noise detector. Out of these, 49,023 came from files labelled as containing sperm whale clicks and 63,846 came from files labelled as not containing sperm whale clicks. The remaining 49,810 short audio segments were randomly sampled from the long audio files to incorporate a representation of background noise. It is important to note that these latter segments were sampled independently from the impulsive noise detections.
4.3 Feature extraction methods and training of 4-min audio file classifier
@@
A total of 28 TCN models with optimised parameters and structure (Table 6) were trained on the collated 4-minute audio dataset for different forms of feature extraction (Figure 4). The best performance both in terms of high recall and low FPR is obtained using a 2D-VAE on spectrogram representations, followed by multi-channel RMS measures and a 1D-VAE on spectral profile. The performance of TCNs applied to encoded representations of waveforms yields the worst results every time, suggesting that the data inputted into the VAE was too noisy and variable to extract the meaningful information needed for this task.
@@
Table 6: Layers and parameters of the TCN used for recording classification, where $N$ is the number of non-overlapping windows in the file and $m$ the number of extracted features.
@@
Figure 5: Box plot displaying recall (orange) and FPR (grey) scores on validation data divided by deployment for the best performing model for 4 different feature extraction types. Dots displaying data for the 9 deployments (Table1) are displayed next to each box plot. The best models will maximize recall while minimizing FPR scores, and thus will have the grey and orange boxes clearly separated. The figure shows that the detector that manages to perform better across data from different deployments is the one that works on sequences of embeddings of 2D spectrograms.
When looking at performance variability across sources, we see that the TCN trained on 2D-VAE embeddings from spectrogram data also sees the highest lowest Recall and the lowest highest FPR (Figure 5). This suggests that this feature extraction method not only performs better overall but also shows a more stable performance across data from different sources.
4.4 The effect of annotation length
On 30-second recordings, the results demonstrate a notable improvement in both higher recalls and lower false positive rates for TCNs trained on most feature extraction methods when compared to the same methods trained on 4-minute recordings (Figure 6, Figure 7). The only two exceptions to this trend are TCNs trained on VAE embeddings from waveforms, which exhibit no improvement and are consistently worse-performing method, and TCNs trained on VAE embeddings from spectrograms, which only show a marginal performance improvement. Notably, for 30-second recording classification, TCNs trained on VAE embeddings for spectral profile show the most improvement and achieve the best overall performance across all tested methods for feature extraction.
5 Discussion
In this study, we developed a novel workflow designed to effectively utilise weakly labelled acoustic datasets, a prevalent challenge in bioacoustics research. Our proposed approach consisted of two stages: an unsupervised feature extraction step for data representation and dimensionality reduction and a classification at the recording level using Temporal Convolutional Networks (TCNs), which incorporated temporal context into the decision making. This approach diverges from conventional segmentation strategies, which risk overlooking crucial contextual cues inherent in the longer intervals at which weak labels are set. We found that with the proposed approach TCNs were able to classify recordings of 4-minutes with a performance comparable to those seen in human analysts [66].
We explored the effectiveness of both handpicked acoustic indices and Variational Autoencoders (VAEs) for unsupervised extraction of features. Results showed that both handpicked features and VAE embeddings were able to provide enough information for a TCN to obtain recall rates of over 0.7 while maintaining a false detection rate below 0.16. TCNs trained on VAE embeddings of spectrogram representations performed marginally better on the task of classifying 4-minute recordings. However, the results showed that handpicked acoustic features were significantly more prone to overfitting to some datasets and their performance varied greatly by deployment, while VAE embeddings achieved a more stable performance across data source. This finding is particularly significant to the bioacoustics community as it highlights the capacity of VAEs to offer a consistent and reliable feature representation, regardless of variations in data sources and deployment conditions.
The study also revealed that classifiers trained and tested on shorter duration recordings (30 seconds instead of 4 minutes) consistently exhibited superior performance. While this outcome was anticipated, one must remember that the shorter dataset comprised recordings from only two sources, potentially contributing to lower data variation. However, what is worth noting is that in this scenario TCNs trained on embeddings of spectral profiles outperformed those trained on the embeddings of 2D spectrograms, which were the best performing method for the 4-minute audio files. This suggests that the length of the final vector of acoustic features became a limiting factor in the classifier’s performance for 4-minute recordings. Consequently, methods generating longer sequences yielded better performances, even at the expense of time resolution. However, for classification at shorter time scales, these features extracted at a higher resolution significantly benefit the TCN step in the classification task.
6 Conclusion
Our study suggests a way forward in leveraging numerous existing annotated bioacoustics datasets to train automatic classification models, effectively overcoming previous limitations associated with weak labels. Notably, our ability to train the model on a scale as large as 4 minutes means that marine scientists can readily apply this approach to nearly all available annotated data without additional auditing, enabling for the processing of vast amounts of data efficiently. This capability opens doors for researchers to analyze higher volumes of marine acoustic data. Beyond marine science, our proposed approach has potential applicability in other acoustic fields, including soundscape analysis, environmental monitoring, or healthcare. Future research should focus on validating the generalisability of the best-performing models on data from new sources.
Acknowledgements
We would like to express our gratitude to Dr Kirsten Young, Thomas Webber, Asociación TURSIOPS, Dr Felicia Vachon, Dr Hal Whitehead, Dr Paul J. Wensveen, Michelle Dutro, Caroline Haas, the Marine Conservation Research/International Fund for Animal Welfare, and Greenpeace International for supplying the essential data for this research. Collection of the Iceland data set was made possible by the Rannís Icelandic Research Fund (number 207081).
Author declarations
Conflicts of Interest
The authors have no conflicts to disclose.
Data availability
7 Supplementary material
Algorithm 1 Impulsive noise detector
Input:
           $w$: waveform vector
Output:
           Start and end indices of detected impulsive noises (clicks)
$\alpha_{n1}$ = 0.00001 $\triangleright$ Long filter 1 (within impulsive noise detection)
$\alpha_{n2}$ = 0.000001 $\triangleright$ Long filter 1 (outwith impulsive noise detection)
3:$\alpha_{s}$ = 0.01 $\triangleright$ Short filter
4:$threshold=6dB$triangleright Minimum SNR to trigger detection
5:$\alpha_{n}$ = $\alpha_{n1}$triangleright Set initial values for signal and noise
6:S = $|w_{0}|$triangleright Set initial signal value as absolute value of the first sample of the waveform
7:N = $Median(|w|)$triangleright Set initial noise value as median for the absolute value of the waveform
8:detection=0 $\triangleright$ This variable stores information on whether a detection is active
9:for $i=256,\ldots length(w)$ do $\triangleright$ Ignore first 256 samples of waveform
10:     if detection then $\triangleright$ Choose long filter
11:         $\alpha_{n}$ = $\alpha_{n1}$
12:     else
13:         $\alpha_{n}$ = $\alpha_{n2}$
14:     end if
15:     $N=\alpha_{n}\times|w_{i}|+(1-\alpha_{n})\times N$triangleright Compute noise and signal levels
16:     $S=\alpha_{s}\times|w_{i}|+(1-\alpha_{s})\times S$
17:     $SNR=20\times\log_{10}(S)-20\times\log_{10}(N)$triangleright Compute SNR
18:     if $SNR>threshold$ then
19:         detection = 1
20:         Store $i$ if it is first index of the detection
21:     else
22:         detection=0
23:         Store $i-1$ if $i$ is first index out of detection
24:     end if
25:end for

### v2

@@ II.2 Pre-processing
An impulsive noise detector algorithm was used to identify and select up to 10 transient sounds with the highest amplitude within each 4 minute segment, regardless of what the recording had been labeled as. Transient sounds were detected using a single pole filter with a 6dB threshold (see Alg. 1 shown in FIG.  2). Subsequently, waveform segments of varying lengths were extracted from these transients to train different feature engineering methods, ensuring the random placement of amplitude peaks within each segment. Additionally,5 randomly sampled segments per file where no transients was detected were included in order to incorporate examples of background noise into the optimisation of feature engineering methods. The proposed approach acknowledges that, while impulsive noises detected in a 4-min sequence might not necessarily share the label of the parent audio file, a considerable proportion are expected to do so, such that this procedure would produce a dataset representative of the different sound sources present in the data. In summary, for each 4-min labelled segment, we extract up to 10 top high amplitude sequences plus 5 randomly selected quiet periods to train the feature extractors.
Figure 2: Impulsive noise detector algorithm
For audio classification at the file level, recordings were divided into shorter, non-overlapping windows of two sizes: 512 and 2048 samples (0.01s and 0.04s, respectively). The two window lengths were used to investigate the trade-off between temporal resolution and sequence length. This process generated 22,500 and 5,625 windows, respectively, for 4-minute files, and 2,812 and 703 windows for 30-second files (Table 7). The feature vectors detailed in the previous section were concatenated, transforming each audio file into a sequence of size ($m$, $n$), where $m$ represents the number of extracted features and $n$ denotes the number of non-overlapping windows in the file. Each of the $m$ features, both from handcrafted features and VAE embeddings, was normalised using median and standard deviations of the feature on the values obtained on the short audio segment dataset.
II.3 Feature engineering
Features were generated two-fold: (i) expert-based or handcrafted feature selection and (ii) feature extraction via variational autoencoders (VAEs). Whilst the former in more transparent and rooted in marine science [6], it may not necessarily be optimised for the classifier. Feature extraction is not explainable to a human but tends to learn compressed, probabilistic latent representations of the PAM segments, capturing key temporal dynamics and structure in the click trains.
II.3.1 Handcrafted feature generation
@@
Several acoustic features were selected from the non-overlapping short audio segments generated as described in Subsection II.2, namely the root-mean-square of the time-series for a specific frequency band $f_{1},f_{2}]$ ($RMS_{f_{1},f_{2}}$), peak and mean frequencies, spectral width, and energy sum over three frequency bands (1kHz - 4kHz, 4kHz-8kHz, 8kHz-16kHz) (see Figure 3 for an example and Table 2 for formal definition of each handcrafted feature). The choice of these parameters was motivated by earlier works on the detection-classification of odontocete vocalisations [6], and the frequency bands were chosen to correspond to the distinct bandwidths in the sperm whale click spectral profile [17].
@@
Table 2: Handcrafted features
@@
Peak frequency $pkf=argmax\left(s(f)\right)$ $f$: vector of frequencies returned by fft
@@ 5 $RMS_{1-2kHz}$, $RMS_{2-4kHz}$, $RMS_{4-8kHz}$, $RMS_{8-16kHz}$, $RMS_{16-20kHz}$,
Spectral features 7 $RMS_{1-20kHz}$, pkfpkf, $\overline{f}$, $\omega$, $E_{1-4kHz}$, $E_{4-8kHz}$, $E_{8-16kHz}$
II.3.2 Feature Extraction with Variational AutoEncoders (VAE)
VAEs were trained on three different forms of input data: waveforms, spectral profiles, and spectrogram representations. To generate the spectrograms and spectral profiles we used a 512-point Hann window fft with a 50% overlap. Spectral profiles were computed on waveforms of length 512 samples, and frequencies below 1kHz or above 20kHz were discarded, leading to a final sequence of length 200, which was then smoothed using an 8-point moving average. Additionally, longer spectrograms were computed for waveform sections of length 32,768 samples (0.68 seconds), so as the resulting spectrogram to be 128 FTTs in length giving an image of width 128 pixels. Again, frequencies below 1kHz and above 20kHz were discarded to center sperm whale clicks in the spectrogram [17], and the final representation was resized to be of size $\times128$ pixels.
A 2D-VAE that utilizes a ResNet-18 architecture [43], a network that has consistently exhibited good performance in bioacoustics call detection literature [2, 21, 28, 41] was tuned for extracting features from spectograms. On the other hand, waveforms and smoothed spectral profiles both were autoencoded using a 1D-VAE [29]. Each model was trained for different latent sizes: 24, 32, 48 and 64. The 2-D VAE was implemented for two additional latent sizes, 96 and 128. Since VAEs return the parameters of a normal distribution in the latent space, for each input data point encoded into a space of dimension $n$, a vector of size 2nn is outputted where the first half of the vector corresponds to the means of the normal distribution and the second half to the variances. This strategy of mapping data to normal distribution parameters facilitates interpolation between data points, enhancing the robustness of embeddings against various forms of noise and domain shifts [39].
II.4 TCN classification
@@
A TCN was used for classifying full recordings from the sequence of extracted features based on the presence or absence of sperm whale vocalisations. The architecture shown in Figure 4 was adopted due to its proven ability to capture long-term dependencies and timing patterns in acoustic signals [25], which would be key for detecting groups of sperm whales of varying sizes and distinguishing them from other sources of impulsive noise. The TCN was of length $n$ and had $m$ channels.
All TCNs were trained in a supervised manner using the Adam optimizer and negative log likelihood loss function. Hyperparameters, including the number of hidden layers, kernel size, dropout rate, learning rate, and number of levels, were optimized using a grid search strategy (Table 6). The best-performing combination consisted of a batch size of 8, 8 hidden layers of size 25, a kernel size of 20, a dropout rate of 0.4, and a learning rate of 0.001. The final architectures for both the 4-minute and 30-second segment models are summarized in Table 7.
III Results
The performances metrics of the different trained TCNs were compared on an unseen subset of the data reserved for testing. A random +:15 split was used to partition the data into training and testing sets, and that split was consistent across models trained on the same length of recordings. Performance evaluation was conducted based on recall ($TP/P$) and false-positive-rate (FPRFPR; $FP/N$), where $P$ and NN represent the total number of positive and negative files (as annotated), and TPTP and FNFN denote correct and incorrect predictions of the model from the files labelled as containing sperm whales by humans. The best model is one that minimizes FPRFPR while simultaneously maximizing recall.
III.1 Data segments for training
The final 4-minute training set contained 7,668 audio files (Table 4) coming from 8 different deployments, with a 50/50 split between files that were deemed as containing and not containing sperm whale clicks. This dataset included data from different sperm whale populations spanning 20 years, 8 different geographical locations, 7 different recording devices and over 8 other vocalising species that produce echolocation clicks, as well as other sources of impulsive noise. The 30-second dataset contained a total of 9,912 files from 2 different sources, AS and MED (Table 4).
Table 4: The number of recordings in the merged dataset labelled as containing only sperm whales (SW), sperm whales and other sources of impulsive noise, other sources of impulsive noise only, and recordings audited but nothing noted down in the final training set separated by source. For the Balearic data recordings were only marked for the presence of sperm whale vocalisations, so no information on other sound sources was available. The numbers outside of parenthesis refer to the 4-minute dataset, and parenthesis to the 30-second dataset.
@@ III.1 Data segments for training
CAL 173 1 114 60 348
CS 773 560 720 607 2,660
ICE 365 558 379 544 1,846
MED 626 (3,003) 16 (31) 47 (230) 594 (2,796) 1,283 (6,060)
Total 2,595 (4903) 1,143 (57) 1,383 (1285) 2,547 (3667) 7,668 (9912)
@@
A total of 101,210 short segments were identified to train the feature engineering methods to train the feature extraction methods. 71,893 of these were sampled using an impulsive noise detector. Out of these, 37,638 came from recordings labelled as containing sperm whale clicks and 34,206 came from recordings labelled as not containing sperm whale clicks. The remaining 29,317 short audio segments were randomly sampled from the long audio segments to incorporate a representation of background noise. It is important to note that these latter segments were sampled independently from the impulsive noise detections.
III.2 Feature extraction methods and training of 4-min audio file classifier
@@
A total of 24 TCN models with optimised parameters and structure (Table 5) were trained on the collated 4-minute audio dataset for different forms of feature extraction (Figure 5). The best performance both in terms of high recall and low FPR is obtained using a 2D-VAE on spectrogram representations, followed by multi-channel RMS measures and a 1D-VAE on spectral profile. The performance of TCNs applied to encoded representations of waveforms yields the worst results every time, suggesting that the data inputted into the VAE was too noisy and variable to extract the meaningful information needed for this task.
Figure 6: Box plot displaying recall (orange) and FPR (grey) scores on validation data divided by deployment for the best performing model for 3 different feature extraction types. Dots displaying data for the 8 deployments (Table1) are displayed next to each box plot. The best models will maximize recall while minimizing FPR scores, and thus will have the grey and orange boxes clearly separated. The figure shows that the detector that manages to perform better across data from different deployments is the one that works on sequences of embeddings of 2D spectrograms.
When looking at performance variability across sources, we see that the TCN trained on 2D-VAE embeddings from spectrogram data also sees the highest lowest Recall and the second lowest highest FPR (Figure 6). This suggests that this feature extraction method not only performs better overall but also shows a more stable performance across data from different sources.
III.3 The effect of annotation length
On 30-second recordings, the results demonstrate a notable improvement in both higher recalls and lower false positive rates for TCNs with different input features when compared to the features from the 4-minute recordings (Figure 7, Figure 8). The only two exceptions to this trend are TCNs trained on VAE embeddings from waveforms, which exhibit no improvement and are consistently worse-performing method, and TCNs trained on VAE embeddings from spectrograms, which only show a marginal performance improvement. Notably, for 30-second recording classification, TCNs trained on VAE embeddings for spectral profile show the most improvement and achieve the best overall performance across all tested methods for feature extraction.
IV Discussion
In this study, we developed a novel interdisciplinary workflow designed to effectively utilise weakly labelled acoustic datasets, a prevalent challenge in bioacoustics research. Our proposed approach consisted of four stages: (i) robust dataset curation from multiple geographical recordings to include various sources of environmental and anthropological noise, (ii) feature generation via expert-based feature selection and unsupervised feature extraction via VAEs, (iii) classification at the recording level using TCNs, which incorporated temporal context into the decision making. This approach diverges from conventional segmentation strategies, which risk overlooking crucial contextual cues inherent in the longer intervals at which weak labels are set. We found that with the proposed approach TCNs were able to classify recordings of 4-minutes with a performance comparable to those carried out by human analysts [14].
We explored the effectiveness of both handpicked acoustic indices and VAEs for unsupervised feature extraction. Results showed that both handpicked features and VAE embeddings were able to provide enough information for a TCN to obtain recall rates of over 0.83 while maintaining a false detection rate below 0.13. TCNs trained on VAE embeddings of spectrogram representations performed marginally better on the task of classifying 4-minute recordings. Our results also showed that handpicked acoustic features were more prone to overfitting to some datasets and their recall rates varied significantly by deployment, while VAE embeddings achieved a more consistent performance across data sources. This finding is particularly significant to the bioacoustics community as it highlights the capacity of VAEs to offer a consistent and reliable feature representation, regardless of variations in data sources and deployment conditions.
Our approach differs from previous sperm whale detection approaches through three key innovations. First, we employ TCNs to model long-range dependencies in acoustic sequences, enabling detection decisions at biologically-relevant timescales for this species (e.g., minute-level click trains) rather than isolated click-level analysis typical of traditional energy detectors [19, 34, 3]. Second, we achieve robust performance over these long timescales without the need for manually set thresholds [31, 49] or finely labeled data [14]. Third, we integrate VAEs as feature extractors, simultaneously improving computational efficiency in the classification step while preserving enough information to distinguish sperm whale vocalizations from ambient noise. This combination of large-scale temporal modeling, reduced annotation dependence, and efficient data compression addresses fundamental limitations in current passive acoustic monitoring pipelines.
To our knowledge, this study also presents the largest and most varied sperm whale acoustic dataset curated to date, combining recordings from six different studies spanning decades, diverse geographic regions, and multiple deployment conditions. This variability, which includes differences in background noise, vocalising species, and recording set ups, is critical to training robust, generalizable models. This dataset provides a reference point for future research on weakly labeled bioacoustic detection, addressing a key need for open, high-quality bioacoustics baseline datasets [13].
Differences in performance between data sources could stem from variations in noise levels and recording conditions [4, 36], but annotation protocols and survey designs may also play a significant role. For example, aurally auditing four minutes of single-channel data may lead to missed detections of fainter, less regular sperm whale clicks, whereas manually reviewing each click in isolation allows for more thorough annotation. Moreover, the purpose of the study that generated a particular dataset also influences the detectability of the animals in the acoustic recordings. For example, the CS dataset was collected during a dedicated sperm whale behavioral study, where a vessel actively tracked whales to study social interactions. Consequently, the recordings predominantly contain high-amplitude, well-defined clicks from nearby individuals, introducing a proximity bias in the vocalization examples. In contrast, datasets from static hydrophones or line surveys are more likely to include faint, distant clicks, which poses challenges for both human annotators and automated detectors. When comparing detector performance across datasets, reported metrics reflect both the algorithm’s efficacy and dataset-specific annotation biases. All results must therefore be interpreted within their recording and annotation context.
The present study also revealed that classifiers trained and tested on shorter duration recordings (30 seconds instead of 4 minutes) consistently exhibited superior performance. Although this outcome was anticipated, it should be noted that the shorter dataset comprised recordings from only two sources, potentially contributing to lower data variation. However, what is worth noting is that in this scenario TCNs trained on embeddings of spectral profiles outperformed those trained on the embeddings of 2D spectrograms, which were the best performing method for the 4-minute audio files. This suggests that the length of the final vector of acoustic features became a limiting factor in the classifier’s performance for 4-minute recordings. Consequently, methods that generate shorter sequences yielded better performances, even at the expense of time resolution. However, for classification at shorter time scales, these features extracted at a higher resolution significantly benefit the TCN step in the classification task.
V Conclusion
Our study suggests a way forward in leveraging numerous existing annotated bioacoustics datasets to train automatic classification models, effectively overcoming previous limitations associated with weak labels. Notably, our ability to train the model on a scale as large as 4 minutes means that marine scientists can readily apply this approach to nearly all available annotated data without additional auditing, enabling for the processing of vast amounts of data efficiently. This capability opens doors for researchers to analyze higher volumes of marine acoustic data. Beyond marine science, our proposed approach has potential applicability in other acoustic fields, including soundscape analysis, environmental monitoring, or healthcare. Future research should focus on transferability of the best-performing models on distinct unseen new sources. Additionally, research will be carried out on more fine-tuned hybrid human-AI feature selection [23] for improved feature learning, robustness and transparency.
Acknowledgements.
This work was funded by the Scottish Universities Partnership for Environmental Research (SUPER) Doctoral Training Partnership (DTP), between the Universities of St Andrews and Strathclyde. We would like to express our gratitude to Dr Kirsten Young, Thomas Webber, Asociación TURSIOPS, Dr Felicia Vachon, Dr Hal Whitehead, Dr Paul J. Wensveen, Michelle Dutro, Caroline Haas, the Marine Conservation Research/International Fund for Animal Welfare, and Greenpeace International for supplying the essential data for this research. Collection of the Iceland data set was made possible by the Rannís Icelandic Research Fund (number 207081).
AUTHOR DECLARATIONS
Conflicts of Interest
The authors have no conflicts to disclose.
Data availability
Supplementary materials
@@ Acknowledgements.
Table 5: Layers and parameters of the TCN used for recording classification, where NN is the number of non-overlapping windows in the file and $m$ the number of extracted features.
@@
RMS 1, 3, 5 512 samples 22500 2812
2048 samples 5625 703
Spectral parameters 7 512 samples 22500 2812
2048 samples 5625 703
1D-VAE on waveform 8, 16, 24, 32 512 samples 22500 2812
2048 samples 5625 703
1D-VAE on spectral profile 8, 16, 24, 32 512 samples 22500 2812
2D-VAE on spectrogram 8, 16, 24, 32, 64, 128 32768 samples 352 44